OpenStudy (anonymous):

Solve for x x/15 = 3 4/5

OpenStudy (alexandervonhumboldt2):

first convert to improper fraction: a b/c=(a*c+b)/c 3 4/5=(3*5+4)/5=19/5

OpenStudy (alexandervonhumboldt2):

now we have x/15=19/5

OpenStudy (alexandervonhumboldt2):

mutliply : x=19/5*15

OpenStudy (alexandervonhumboldt2):

x=19*3=???

OpenStudy (anonymous):